OpenStudy (anonymous):

Given that one root of the equation x^2 + 2px - q = o is twice the other, express q in terms of p.

OpenStudy (amistre64):

you recall the quadratic formula?

OpenStudy (anonymous):

We were told not to do that method as it is extremely long and time consuming

OpenStudy (anonymous):

Can you show me the sum of roots and product of roots method of doing it?

sam (.sam.):

can we use teh sum and product of roots?

OpenStudy (anonymous):

i think so...

OpenStudy (amistre64):

im do not recall hearing of those kinds of methods. i do not see how the quad formula is long and consuming tho .... \[\frac{-b-\sqrt{b^2-4ac}}{a}=\frac{-b+\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Alright, sure thing i will try to understand the steps.

OpenStudy (amistre64):

in any case, the roots will still be a result of the quad formula .... not too sure how you would be able to define them otherwise

OpenStudy (anonymous):

Can you show me the steps please?

OpenStudy (amistre64):

\[\frac{-b-\sqrt{b^2-4ac}}{a}=\frac{-b+\sqrt{b^2-4ac}}{2a}\] \[-2b-2\sqrt{b^2-4ac}=-b+\sqrt{b^2-4ac}\] \[-2(2p)-2\sqrt{(2p)^2+4q}=-(2p)+\sqrt{(2p)^2+4q}\] \[-2p=3\sqrt{(2p)^2+4q}\] \[-\frac23p=\sqrt{4p^2+4q}\] \[-\frac23p=2\sqrt{p^2+q}\] \[-\frac13p=\sqrt{p^2+q}\] \[\frac19p^2=p^2+q\] \[-\frac89p^2=q\] maybe

OpenStudy (anonymous):

I see so this is how i do it. Thank you :D *Thumbs up*

OpenStudy (anonymous):

Thanks dude.

OpenStudy (amistre64):

this is how i would approach it, yes.

OpenStudy (anonymous):

Can i ask you to help me solve a similar question?

OpenStudy (amistre64):

dunno, i only have so many good ideas in a day ...

OpenStudy (anonymous):

in 3 mins time

OpenStudy (anonymous):

Can i swop the denominator ?

OpenStudy (anonymous):

Will it affect the end result?

OpenStudy (amistre64):

what do you mean be "swap the denominator"?

OpenStudy (anonymous):

the a and 2a swop the position

OpenStudy (anonymous):

because gen formula is -b positive/negative square root (b square -4ac) over 2a

OpenStudy (anonymous):

The denominator represent the roots and a is the unknown variable right?

OpenStudy (amistre64):

the quadratic formula has 2 parts to it, a vertex axis called a discriminant, and a +- part that measures an equal distance to both sides of that axis to define the zeros: r1 <= vertex <= r2 \[\frac{-b}{2a}-\frac{\sqrt{b^2-4ac}}{2a}~\le~\frac{-b}{2a}~\le~\frac{-b}{2a}+\frac{\sqrt{b^2-4ac}}{2a}\]

OpenStudy (amistre64):

the phrase "2 times as large" suggests to me that the left side has to double to equal the right side

OpenStudy (anonymous):

what if it's like the root differ by 2

OpenStudy (anonymous):

or you need to see the whole question

OpenStudy (amistre64):

the left and right sides of the middle are the roots; if we keep it clean and just call them r1 and r2 |r1 - r2| = 2 expresses the difference between them as "2"

OpenStudy (anonymous):

I see wow this logic is quite thought provoking..

OpenStudy (anonymous):

okay moving on to the next question now